OpenStudy (wwe123):

given that α and β are the roots of the equation x²= x+7

OpenStudy (anonymous):

0 = x^2 - x - 7 0 = x(x - 1) - 7

OpenStudy (wwe123):

prove that \[(a) \frac{ 1 }{\alpha} =\frac{\alpha-1}{ 7}\]

OpenStudy (anonymous):

\[x^2 - x - 7 = 0\]

OpenStudy (anonymous):

go ahead an solve, see what you get

OpenStudy (anonymous):

6 and 7, but that doesn't work in a proof...

OpenStudy (anonymous):

ooops 7 and 8

hartnn (hartnn):

since a is the root, u can write a^2 = a+7 a^2-a=7 a(a-1)=7 a-1 = 7/a 1/a = (a-1)/7

OpenStudy (anonymous):

a=alpha

hartnn (hartnn):

a(a-1)=7 a-1 = 7/a<-------dividing both sides by 7 1/a = (a-1)/7<------dividing both sides by (a-1)

OpenStudy (wwe123):

yes

OpenStudy (anonymous):

He just used factor theorem.

OpenStudy (anonymous):

U know factor theorem?

OpenStudy (anonymous):

neither 7 nor 8 area solutions

OpenStudy (wwe123):

plz solve part (b) \[\alpha^{3} =8\alpha+7\]

hartnn (hartnn):

a^2 = a+7 a^3 = a^2 + 7a =a+7+7a =8a+7

OpenStudy (anonymous):

^ jobs done

OpenStudy (wwe123):

thanks

hartnn (hartnn):

did u understand ?